triangle A is similar to triangle B. side length of triangle A is 3 inches .side length of triangle B is 12 inches. the area of A is 10. what is the area of B ?
a 20 b 40 c 160 d 640 e 40640

Answer by drk(1908)

Based on the information, we can write the ratio of side A to side B as
3/12 = 1/4.
To go from side ratio to area ratio, square the side ratio.
(1/4)^2 = 1/16. This means that triangle B has an area 16 times bigger than the area of triangle A. Since the area of triangle A is 10, 10 * 16 = 160. This is choice [ C ].